# Following the Autoregressive Nature of LLM Embeddings via Compression and Alignment

### v2

@@ 1 Introduction
Text embeddings, which represent the semantic content of natural language text as vectors, are extensively utilized in domains such as information retrieval, semantic similarity assessment, retrieval-augmented generation (RAG) (Khandelwal et al., 2020; Shi et al., 2024; Deng et al., 2023; Ding et al., 2024; Xu et al., 2024), and data attribution (Beigi et al., 2024). Traditional text embedding models typically employ transformer-based architectures with encoder-only designs, including examples like Bert (Devlin et al., 2019), DeBERTa (He et al., 2021) and MPNet (Song et al., 2020), and are trained using contrastive learning.
@@
After extensive pre-training on a large-scale corpus, LLMs have outperformed previous encoder-only small models (Deng et al., 2022) and demonstrated strong adaptability across diverse downstream tasks (Zhao and Zhang, 2024; Wang et al., 2024b; Zhao et al., 2025). Recently, contrastive learning has been directly applied to decoder-only LLMs, which are trained to generate embedding vectors based on task-specific instructions, enabling adaptability to various embedding scenarios Lee et al. (2024); BehnamGhader et al. (2024). Despite initial advancements, training a high-performance 7B-scale text embedding model using this approach remains highly resource-intensive. It typically requires millions of triplets Wang et al. (2024a); Li et al. (2024b, 2023) and substantial computational power, including thousands of hours on an A100 80GB GPUs Muennighoff et al. (2024); Ma et al. (2024), even with the application of Parameter-Efficient Fine-Tuning (PEFT) Hu et al. (2022); Dao (2024). The high resource consumption might reasonably be attributed to the inability of the discriminative contrastive learning method to fully harness the capabilities of generative LLMs Li et al. (2024a). Firstly, the constraint of unidirectional attention in LLMs leads to the aggregation of information in the hidden state of the output layer corresponding to the final token. However, as LLMs are optimized for next-token prediction, this hidden state can only represent the semantics of the next token (local) rather than the semantics of the input text itself (global). Consequently, employing this hidden state directly in contrastive learning requires additional training time and computational cost to transition from a localized to a more global semantic representation. Secondly, the hidden state in LLMs is used to generate the probability distribution of the next token, whereas contrastive learning optimizes the cosine distance between the hidden states of different texts. This divergence in optimization objectives introduces additional training costs. This raises an important question: Is it feasible to develop a method that follows the auto-regressive nature while generating high-quality text embeddings and significantly reducing resource requirements?
@@
Experimental results on 10 STS datasets demonstrate that AutoRegEmbed outperforms traditional contrastive learning methods while utilizing the same computational resources, making it a highly efficient and scalable solution. Remarkably, even with a limited number of training samples, AutoRegEmbed achieves performance on par with the current state-of-the-art (SOTA) models, showcasing its superior ability to learn robust and generalizable representations from scarce data. As shown in Figure 1, the Pareto frontier of AutoRegEmbed consistently outperforms traditional contrastive learning methods, demonstrating a more optimal trade-off between computational efficiency and performance. This indicates that AutoRegEmbed achieves superior representation learning while maintaining a balanced resource utilization.
@@ 2 Related Works
Early approaches include SentenceBERT Reimers and Gurevych (2019) (supervised) and SimCSE Gao et al. (2021) (unsupervised), which leverage contrastive learning to generate high-quality text embeddings using small encoder-only models. Another area of focus has been improving the isotropy of embedding spaces. Works such as BERT-flow Li et al. (2020) (flow models) and BERT-whitening Su et al. (2021) (linear transformations) address the anisotropic properties of embeddings. Meanwhile, multi-stage contrastive learning Li et al. (2023); Ni et al. (2022); Wang et al. (2022) has further advanced text embeddings by combining pre-training on large weakly supervised datasets with fine-tuning on smaller high-quality datasets. Inspired by instruction fine-tuning, recent research Su et al. (2023); Asai et al. (2023) has shifted toward using text paired with instructions to enhance the generalization and transferability of text embeddings in complex scenarios. However, the performance of methods based on early models is limited, due to their reliance on models with relatively small parameter counts.
@@
Many studies have focused on transforming LLMs into text embedding models through contrastive learning fine-tuning. RepLLaMA (Ma et al., 2024), for example, follows the DPR (Karpukhin et al., 2020) pipeline, using the hidden state of the last token generated by LLaMA as a text embedding vector and applying contrastive learning fine-tuning. Recognizing that the unidirectional attention mechanism in LLMs may limit text embedding quality, LLM2Vec (BehnamGhader et al., 2024) introduces a bidirectional attention mechanism combined with average pooling to enhance embedding quality. NV-Embed (Lee et al., 2024) takes this further by incorporating an additional Latent Attention Layer to generate pooled embeddings. bge-en-icl Li et al. (2024b) suggests that retaining the original framework of LLMs and leveraging in-context learning is the optimal approach for generating text embeddings. Some studies Wang et al. (2024a) even use synthetic data generated by LLMs, rather than real-world data, for fine-tuning and achieve competitive performance on the MTEB leaderboard Muennighoff et al. (2023). However, these approaches often overlook the fundamental differences between language modeling and contrastive learning, failing to fully leverage the potential of LLMs. More closely related to our work is Llama2Vec Li et al. (2024a), which proposes two pretext tasks to enable unsupervised adaptation of LLMs, followed by contrastive learning fine-tuning to achieve better performance. In contrast, our approach achieves strong results without any need for contrastive learning fine-tuning, as our task fully exploits the inherent potential of LLMs.
@@ 3 Method
Figure 2: Overall framework of AutoRegEmbed. Firstly, we perform the information compression task to inject key information from the context and instruction into the compressed tokens. Then, we optimize the conditional probability distribution of these tokens to align the distributions of $e_{q,I_{\mathrm{next}}}$ and $e_{d^{+},I_{\mathrm{self}}}$ as closely as possible through $S_{1}(q,d^{+})$, while increasing the probability of $e_{q,I_{\mathrm{next}}}$ generating positive samples and reducing the probability of $e_{q,I_{\mathrm{next}}}$ generating negative samples through $S_{2}(d^{+},d^{-};q)$.
@@ 3.2 Information Compression: from Discriminative to Generative Embeddings
In decoder-based LLMs, embeddings are typically generated by extracting the hidden state of the final token in the input sequence. However, this approach primarily captures the semantics of the first output token rather than encoding the global semantics of the entire input. Various pooling techniques, such as average pooling and attention pooling, have been explored to mitigate this limitation, yet they introduce their own challenges. The average pooling method, which computes the mean of all token hidden states, does not necessarily encapsulate global semantics but instead serves as a mechanism for "convexity preservation." Conversely, attention pooling modifies the attention mechanism or introduces additional parameters, thereby altering the original architecture of LLMs. Such modifications deviate from the model’s pre-training design and can lead to unintended consequences, as prior studies Li et al. (2024b) indicate that maintaining the original LLM framework often yields optimal performance. To enable LLMs to generate embeddings that represent global semantics, we introduce an information compression task. This task compels LLMs to reconstruct the original target using a compressed embedding derived from the input text. Given that this compressed embedding models the conditional probability distribution of the target, we designate it as the generative embedding to contrast it with the discriminative embedding produced by conventional pooling approaches.
@@ 3.3 Conditional Distribution Alignment: from Data-Point to Distribution Perspective
After addressing the global semantic representation issue of the embedding vector, we also require the embedding vector to meet the criteria of alignment and uniformity. Existing studies Wang and Isola (2020) have provided specific definitions for these properties. Alignment is typically expressed by the following formula:
$\mathrm{Alignment}(f,\alpha)=\underset{(q,d^{+})\in p_{\mathrm{pos}}}{\mathbb{E}}\parallel f(q)-f(d^{+})\parallel^{\alpha}_{2},$
where $\alpha>0$ is a parameter used to adjust the weight of the distance between positive sample pairs ($q$ and $d^{+}$), and $p_{\mathrm{pos}}(\cdot,\cdot)$ represents the distribution of positive sample pairs. The smaller $\mathrm{Alignment}(f,\alpha)$ is, the better the alignment of the embedding vector generated by $f$. Uniformity measures how evenly the embedding is distributed, commonly expressed by the following formula:
$\mathrm{Uniformity}(f,\alpha)=\mathrm{log}\underset{(q,d)\in p_{\mathrm{data}}}{\mathbb{E}}e^{-t\parallel f(q)-f(d)\parallel^{2}_{2}},$
where $t>0$ and $p_{\mathrm{data}}$ represents the data distribution. In general, we optimize these two properties asymptotically using a contrastive loss, such as InfoNCE,
@@
where $\mathrm{D}(\cdot,\cdot)$ is any function that measures the divergence between two probability distributions. In addition, since the conditional probability distribution of the embedding can be adjusted based on the given instruction, we can compute the logarithmic probability of the text embedding $e_{q,I_{\mathrm{next}}}$ generating positive or negative samples to measure the similarity between the text $q$ and the positive $d^{+}$ and negative $d^{-}$ samples. Here, the instruction $I_{\mathrm{next}}$ is similar to “find documents that can answer this question”, which ensures that the embedding $e_{q,I{\mathrm{next}}}$ generated from text $q$ produces positive samples after passing through the decoder. For positive and negative samples, we use the instruction $I_{\mathrm{self}}$ similar to “find sentences that are semantically similar to this text”, so that their embeddings, $e_{d^{+},I_{\mathrm{self}}}$ and $e_{d^{-},I_{\mathrm{self}}}$, generate themselves after passing through the decoder.
Building on the above insights and incorporating the structure of InfoNCE, we empirically derive the final loss function:
@@
$-\beta\ \mathrm{log}\frac{p_{\theta_{E}}(d^{-}|e_{q,I_{\mathrm{next}}})}{p_{\mathrm{ref}}(d^{-}|e_{q,I_{\mathrm{next}}})}),$
where $\tau$ and $\beta$ are temperature parameters, and $p_{\Theta_{L_{E}}}$ represents the initial model. We use the Sigmoid function $\sigma(\cdot)$ to normalize the similarity measured from the conditional probability distribution to the range [0,1], ensuring maximum consistency with the range of cosine distance. $S_{1}$ represents the similarity function between text $q$ and the positive sample $d^{+}$. We define it by measuring the absolute value of the difference in the logarithmic probability of their corresponding embeddings, $e_{q,I_{\mathrm{next}}}$ and $e_{d^{+},I_{\mathrm{self}}}$, generating the positive sample $d^{+}$. To minimize this difference, we apply the absolute value function. In addition, we then add a negative sign to ensure that the value of $S_{1}$ increases as the similarity between $q$ and $d^{+}$ increases. $S_{2}$ calculates the difference between the logarithmic probabilities of generating positive and negative samples for text $q$, similar to DPO (Rafailov et al., 2023). We amplify this difference to boost the probability of embedding $e_{q,I_{\mathrm{next}}}$ generating positive samples and decrease the probability of generating negative samples. We normalize the probabilities by dividing them by the corresponding values from the initial model to account for the length discrepancy between positive and negative samples.
@@
AutoRegEmbed-LLaMA2 7B 85.50 79.07 79.57 86.90 83.28 88.45 86.57 88.61 66.16 86.59 84.35(7)/83.07(10) 50,000(16,382)
AutoRegEmbed-Mistral 7B 86.69 80.21 78.33 86.22 82.36 88.42 86.43 88.70 64.27 87.05 84.15(7)/82.87(10) 50,000(16,382)
AutoRegEmbed-LLaMA2 7B 85.62 81.93 78.84 86.76 84.01 89.43 87.72 89.04 66.77 87.96 85.24(7)/83.81(10) 274,951(16,382)
Table 1: Results on STS tasks (Spearman correlation scaled by 100x). The parentheses in the Avg. column indicate the number of datasets used to compute the average. Vol. denotes the number of training triplets, while the numbers in brackets indicate the instruction fine-tuning data used by AutoRegEmbed during the information compression stage. The symbol “~” denotes an estimated value. " " represents our own fair baselines, and we apply a grid search to ensure optimal performance. "-L" and "-M" denote the hidden state of the last token and the average pooling of all token hidden states, respectively. The symbol $\clubsuit$ indicates that not all data are open source.
@@ 4.1 Experimental Settings
Previous studies Gao et al. (2021); Li et al. (2020) highlight that a key goal of text embedding is to cluster semantically similar sentences. Following this approach, we use the MTEB Muennighoff et al. (2023) evaluation framework to evaluate AutoRegEmbed on ten semantic text similarity datasets, including STS12 (Agirre et al., 2012), STS13 (Agirre et al., 2013), STS14 (Agirre et al., 2014), STS15 (Agirre et al., 2015), STS16 (Agirre et al., 2016), STS17 (Cer et al., 2017), STS22 (Chen et al., 2022), STS-B , BIOSSES and SICK-R. Each pair of text in the STS dataset is labeled with a similarity score ranging from 0 to 5 or 0 to 4, indicating their semantic similarity. The evaluation metric is the Spearman correlation between the similarity scores predicted by the model and the scores annotated by humans.
@@
In the information compression stage, we use the training set of the instruction fine-tuning dataset PWC (Ge et al., 2024), which includes a diverse range of instruction types, as the training data. The original dataset contains 241,564 (context, instruction, target) samples. To reduce redundancy caused by repeated contexts, we remove duplicates, resulting in the PWC-Unique dataset with 16,382 samples as the final training data. In the conditional distribution alignment stage, we use NLI data (Wang et al., 2024a) and (Chen et al., 2024) from previous studies as training data. The former contains 50,000 samples, while the latter consists of 274,951 samples. Each sample includes an anchor, a positive sample, and a negative sample. Unless otherwise specified, the AutoRegEmbed results presented in the experiment section are based on training with 50,000 samples.

### v3

@@ 1 Introduction
Text embeddings, which represent the semantic content of natural language text as vectors, are extensively utilized in domains such as information retrieval (Xia et al., 2015), semantic similarity assessment, retrieval-augmented generation (RAG) (Khandelwal et al., 2020; Shi et al., 2024; Deng et al., 2023; Ding et al., 2024; Xu et al., 2024b, 2025), LLMs-based agents (Chen et al., 2024a; Xu et al., 2024a) and data attribution (Beigi et al., 2024). Traditional text embedding models typically employ transformer-based architectures with encoder-only designs, including examples like Bert (Devlin et al., 2019), DeBERTa (He et al., 2021) and MPNet (Song et al., 2020), and are trained using contrastive learning.
@@
After extensive pre-training on a large-scale corpus, LLMs have outperformed previous encoder-only small models (Deng et al., 2022) and demonstrated strong adaptability across diverse downstream tasks (Zhao and Zhang, 2024; Wang et al., 2024b; Zhao et al., 2025; Duan et al., 2025; Liu et al., 2025). Recently, contrastive learning has been directly applied to decoder-only LLMs, which are trained to generate embedding vectors based on task-specific instructions, enabling adaptability to various embedding scenarios Lee et al. (2024); BehnamGhader et al. (2024). Despite initial advancements, training a high-performance 7B-scale text embedding model using this approach remains highly resource-intensive. It typically requires millions of triplets Wang et al. (2024a); Li et al. (2024b, 2023) and substantial computational power, including thousands of hours on an A100 80GB GPUs Muennighoff et al. (2024); Ma et al. (2024), even with the application of Parameter-Efficient Fine-Tuning (PEFT) Hu et al. (2022); Dao (2024). The high resource consumption might reasonably be attributed to the inability of the discriminative contrastive learning method to fully harness the capabilities of generative LLMs Li et al. (2024a). Firstly, the constraint of unidirectional attention in LLMs leads to the aggregation of information in the hidden state of the output layer corresponding to the final token. However, as LLMs are optimized for next-token prediction, this hidden state can only represent the semantics of the next token (local) rather than the semantics of the input text itself (global). Consequently, employing this hidden state directly in contrastive learning requires additional training time and computational cost to transition from a localized to a more global semantic representation. Secondly, the hidden state in LLMs is used to generate the probability distribution of the next token, whereas contrastive learning optimizes the cosine distance between the hidden states of different texts. This divergence in optimization objectives introduces additional training costs. This raises an important question: Is it feasible to develop a method that follows the auto-regressive nature while generating high-quality text embeddings and significantly reducing resource requirements?
@@
Experimental results demonstrate that AutoRegEmbed outperforms traditional contrastive learning methods while utilizing the same computational resources, making it a highly efficient and scalable solution. Remarkably, even with a limited number of training samples, AutoRegEmbed achieves performance on par with the current state-of-the-art (SOTA) models, showcasing its superior ability to learn robust and generalizable representations from scarce data. As shown in Figure 1, the Pareto frontier of AutoRegEmbed consistently outperforms traditional contrastive learning methods, demonstrating a more optimal trade-off between computational efficiency and performance. This indicates that AutoRegEmbed achieves superior representation learning while maintaining a balanced resource utilization.
@@ 2 Related Works
Early approaches include SentenceBERT Reimers and Gurevych (2019) (supervised) and SimCSE Gao et al. (2021) (unsupervised), which leverage contrastive learning to generate high-quality text embeddings using small encoder-only models. Inspired by instruction fine-tuning, recent research Su et al. (2023); Asai et al. (2023) has shifted toward using text paired with instructions to enhance the generalization and transferability of text embeddings in complex scenarios. At this stage, several studies have investigated the use of generative tasks to enhance text embeddings. For example, coCondenser (Gao and Callan, 2022) proposes a pre-training strategy that improves the expressiveness of the [CLS] token using unsupervised corpora. PaSeR (Wu and Zhao, 2022) and RetroMAE (Xiao et al., 2022) adopt encoder-decoder architectures: PaSeR emphasizes the reconstruction of key phrases, while RetroMAE employs a masked-token objective to recover the original sentence. However, these approaches do not incorporate contrastive supervision signals into their generative objectives. Moreover, they are evaluated only on smaller-scale models, making them fundamentally different from our method in both design philosophy and scaling strategy.
@@
Many studies have focused on transforming LLMs into text embedding models through contrastive learning fine-tuning. RepLLaMA (Ma et al., 2024), for example, follows the DPR (Karpukhin et al., 2020) pipeline, using the hidden state of the last token generated by LLaMA as a text embedding vector and applying contrastive learning fine-tuning. Recognizing that the unidirectional attention mechanism in LLMs may limit text embedding quality, LLM2Vec (BehnamGhader et al., 2024) introduces a bidirectional attention mechanism combined with average pooling to enhance embedding quality. NV-Embed (Lee et al., 2024) takes this further by incorporating an additional Latent Attention Layer to generate pooled embeddings. bge-en-icl Li et al. (2024b) suggests that retaining the original framework of LLMs and leveraging in-context learning is the optimal approach for generating text embeddings. Some studies Wang et al. (2024a) even use synthetic data generated by LLMs, rather than real-world data, for fine-tuning and achieve competitive performance on the MTEB leaderboard Muennighoff et al. (2023). However, these approaches often overlook the fundamental differences between language modeling and contrastive learning, failing to fully leverage the potential of LLMs. More closely related to our work is Llama2Vec Li et al. (2024a), which proposes two pretext tasks to enable unsupervised adaptation of LLMs, followed by contrastive learning fine-tuning to achieve better performance. In contrast, our approach achieves strong results without any need for traditional contrastive learning fine-tuning (cosine-based), as our task fully exploits the inherent potential of LLMs.
@@ 3 Method
Figure 2: Overall framework of AutoRegEmbed. Firstly, we perform the information compression task to inject key information from the context and instruction into the compressed tokens. Then, we optimize the conditional probability distribution of these tokens to align the distributions of $e_{q,I_{\mathrm{next}}}$ and $e_{d^{+},I_{\mathrm{self}}}$ as closely as possible through $S_{1}(q,d^{+})$, while increasing the probability of $e_{q,I_{\mathrm{next}}}$ generating positive samples and reducing the probability of $e_{q,I_{\mathrm{next}}}$ generating negative samples through $S_{2}(d^{+},d^{-};q)$. Encoder and decoder share a structure.
@@ 3.2 Information Compression: from Discriminative to Generative Embeddings
In decoder-based LLMs, embeddings are typically generated by extracting the hidden state of the final token in the input sequence. However, this approach primarily captures the semantics of the first output token rather than encoding the global semantics of the entire input. Various pooling techniques, such as average pooling and attention pooling, have been explored to mitigate this limitation, yet they introduce their own challenges. The average pooling method, which computes the mean of all token hidden states, does not necessarily encapsulate global semantics but instead serves as a mechanism for "convexity preservation" Li et al. (2020). Conversely, attention pooling modifies the attention mechanism or introduces additional parameters, thereby altering the original architecture of LLMs. Such modifications deviate from the model’s pre-training design and can lead to unintended consequences, as prior studies Li et al. (2024b) indicate that maintaining the original LLM framework often yields optimal performance. To enable LLMs to generate embeddings that represent global semantics, we introduce an information compression task. This task compels LLMs to reconstruct the original target using a compressed embedding derived from the input text. Given that this compressed embedding models the conditional probability distribution of the target, we designate it as the generative embedding to contrast it with the discriminative embedding produced by conventional pooling approaches.
@@ 3.3 Conditional Distribution Alignment: from Data-Point to Distribution Perspective
After addressing the global semantic representation issue of the embedding vector, we also require the embedding vector to meet the criteria of alignment and uniformity. In general, we optimize these two properties asymptotically using a contrastive loss, such as InfoNCE,
@@
where $\mathrm{D}(\cdot,\cdot)$ is any function that measures the divergence between two probability distributions. Given that the embedding distribution is instruction-dependent (see Section 3.1), we can adopt multiple approaches to measure similarity between q and $^{+}/dd^{-}$." Our empirical approach uses a basic alignment strategy: aligning the probability distributions of q and d via distinct instructions ($I_{\mathrm{next}}$ and $I_{\mathrm{self}}$; see Appendix D). This increases the probability of q generating $d^{+}$ and decreases that of generating $d^{-}$. As shown in Appendix E, this simple alignment outperforms more complex alternatives. Building on the above insights and incorporating the structure of InfoNCE, we empirically derive the final loss function:
@@
$-\beta\ \mathrm{log}\frac{p_{\theta_{E}}(d^{-}_{i}|e_{q,I_{\mathrm{next}}})}{p_{\mathrm{ref}}(d^{-}_{i}|e_{q,I_{\mathrm{next}}})}),$
where $\tau$ and $\beta$ are temperature parameters, and $p_{\Theta_{E}}$ represents the initial model. $p_{ref}$ denotes the log probability of the model generating the target before conditional distribution alignment. We use the Sigmoid function $\sigma(\cdot)$ (see Appendix C for analysis) to normalize the similarity measured from the conditional probability distribution to the range [0,1]to prevent overflow during the exponential operation. $S_{1}$ represents the similarity function between text $q$ and the positive sample $d^{+}$. We define it by measuring the absolute value of the difference in the logarithmic probability of their corresponding embeddings, $e_{q,I_{\mathrm{next}}}$ and $e_{d^{+},I_{\mathrm{self}}}$, generating the positive sample $d^{+}$. To minimize this difference, we apply the absolute value function. In addition, we then add a negative sign to ensure that the value of $S_{1}$ increases as the similarity between $q$ and $d^{+}$ increases. $S_{2}$ calculates the difference between the logarithmic probabilities of generating positive and negative samples for text $q$, similar to DPO (Rafailov et al., 2023). We amplify this difference to boost the probability of embedding $e_{q,I_{\mathrm{next}}}$ generating positive samples and decrease the probability of generating negative samples. We normalize the probabilities by dividing them by the corresponding values from the initial model to account for the length discrepancy between positive and negative samples. Note that during inference, the model still computes similarity between $e_{q,I_{\mathrm{next}}}$ and $e_{d,I_{\mathrm{self}}}$ using inner product matching, enabling seamless integration with existing embedding systems.
@@
AutoRegEmbed-LLaMA2 7B 84.65 81.46 79.98 86.35 83.33 89.21 86.91 87.67 65.90 86.98 84.89(7)/83.24(10) 50,000(16,382)
AutoRegEmbed-Mistral 7B 86.84 80.32 78.92 86.18 83.29 88.98 86.75 88.77 64.53 87.24 $84.53_{(7)}$/83.18(10) 50,000(16,382)
AutoRegEmbed-LLaMA2 7B 85.62 83.87 79.60 87.36 84.29 89.43 87.72 89.46 67.78 87.96 85.75(7)/84.31(10) 274,951(16,382)
AutoRegEmbed-Mistral 7B 87.48 83.90 79.56 87.64 84.11 89.58 87.46 89.87 67.77 88.48 85.82(7)/84.59(10) 274,951(16,382)
Table 1: Results on STS tasks (Spearman correlation scaled by 100x). The parentheses in the Avg. column indicate the number of datasets used to compute the average. Vol. denotes the number of training triplets, while the numbers in brackets indicate the instruction fine-tuning data used by AutoRegEmbed during the information compression stage. The symbol “~” denotes an estimated value. " " represents our own fair baselines, and we apply a grid search to ensure optimal performance. "-L" and "-M" denote the hidden state of the last token and the average pooling of all token hidden states, respectively. The symbol $\clubsuit$ indicates that not all data are open source. Bold indicates the best result, and underline indicates the second-best (suboptimal) result.
@@ 4.1 Experimental Settings
Previous studies Gao et al. (2021); Li et al. (2020) highlight that a key goal of text embedding is to cluster semantically similar sentences. Following this approach, we use the MTEB Muennighoff et al. (2023) evaluation framework to evaluate AutoRegEmbed on ten semantic text similarity datasets, including STS12 (Agirre et al., 2012), STS13 (Agirre et al., 2013), STS14 (Agirre et al., 2014), STS15 (Agirre et al., 2015), STS16 (Agirre et al., 2016), STS17 (Cer et al., 2017), STS22 (Chen et al., 2022), STS-B , BIOSSES and SICK-R. Each pair of text in the STS dataset is labeled with a similarity score ranging from 0 to 5 or 0 to 4, indicating their semantic similarity. The evaluation metric is the Spearman correlation between the similarity scores predicted by the model and the scores annotated by humans. In addition, to evaluate retrieval performance, we assess the model on MS MARCO (Nguyen et al., 2016), NFCorpus (Boteva et al., 2016), and SCIDOCS (Cohan et al., 2020) using nDCG@10 as the evaluation metric.
@@
In the information compression stage, we use the training set of the instruction fine-tuning dataset PWC (Ge et al., 2024), which includes a diverse range of instruction types, as the training data. The original dataset contains 241,564 (context, instruction, target) samples. To reduce redundancy caused by repeated contexts, we remove duplicates, resulting in the PWC-Unique dataset with 16,382 samples as the final training data. In the conditional distribution alignment stage, we use the NLI part of the MEDI (Wang et al., 2024a) and BGE (Chen et al., 2024b) datasets as training data. The former contains 50,000 samples, while the latter consists of 274,951 samples. Each sample includes an anchor, a positive sample, and a negative sample. Unless otherwise specified, the AutoRegEmbed results presented in the experiment section are based on the NLI part of the MEDI. For the retrieval task, we train our model on the MS MARCO training set. Because of the asymmetric nature of the retrieval task, the gap between the cosine distance–based evaluation metric and the conditional probability–based optimization objective becomes pronounced. To mitigate this, we perform an additional epoch of contrastive fine-tuning to better align with the evaluation process. Based on prior studies and empirical observations, the quality of negative samples plays a crucial role in training effectiveness. To enhance negative sampling, we adopt NV-Embed for hard negative mining. Specifically, we randomly select 7 hard negatives from the ranked list positions 30 to 210, treating these as challenging negative samples during training.